# Mapping Dark Matter in the Milky Way using Normalizing Flows and Gaia DR3

### v2

@@ 1 Introduction
Assuming that the phase space density of the stars is in equilibrium $\partial f/\partial t=0$, the 3D acceleration field $-\partial\Phi/\partial x_{i}$ can be derived from knowledge of the stellar phase space density $f$ today. A further derivative of $\Phi$ then gives the total mass density $\rho$; the dark matter density can then be calculated assuming knowledge of the baryonic components.
@@ 2 Gaia DR3
In this analysis, we do not use other spectral information (such as $BP-RP$ color).
Figure 1: Schematic representation of the Solar location (red dot) relative to the Galactic Center (black dot). The 4 kpc observation volume is shown as a transparent grey sphere. The three coordinate systems used in this work are shown: the Galactocentric Cartesian coordinates $(x,y,z)$, the spherical coordinates $(r,\theta,\phi)$, and the cylindrical coordinates $(R,\phi,z)$. The lines through the observational volume with low dust extinction along which we measure accelerations and mass densities are shown in color. In orange, we show two lines at $z=+1.5$ kpc, one varying $r$ and another $\phi$. The two corresponding lines at $z=-1.5$ kpc are shown in purple. The line parameterized by polar arclength $s=r_{\odot}\times(\pi/2-\theta)$ passing through the Solar location is shown in green.
@@
    Our Cartesian coordinate system places the $x-y$ plane in the Galactic disk, with $x=0$ at the Galactic center and the Sun along the $+x$ axis. The $+z$ axis (perpendicular to the disk) is oriented so that the net rotation of the disk stars in the $-y$ direction.
@@
    The cylindrical coordinate system $(R,\phi,z)$ uses the same $z$ axis as the Galactocentric Cartesian coordinates. The Sun is located at $\phi=0$, with positive $\phi$ increasing towards the $+y$ axis.
@@
    The spherical coordinate system $(r,\theta,\phi)$ has the same $\phi$ angle as the cylindrical coordinates, and measures $+\theta$ relative to $+z$ axis, with the Galactic disk at $\theta=\pi/2$. We define the polar arc length above or below the Galactic disk at the Solar radius (along the $\theta$ direction) as $s\equiv r_{\odot}\times(\pi/2-\theta)$, with $r_{\odot}=8.122$ kpc.
From the fundamental kinematic properties (parallax, angular position, etc.) measured by Gaia, the positions and velocities in these three coordinate systems can be obtained, using the parallax to calculate distance and assuming a Galactocentric Solar position and velocity of (8.122, 0.0, 0.0208)8.1220.00.0208(8.122,\,0.0,\,0.0208)( 8.122 , 0.0 , 0.0208 ) kpc [29, 30] and (12.9,245.6, 7.78)12.9245.67.78(-12.9,\,-245.6,\,7.78)( - 12.9 , - 245.6 , 7.78 ) km/s [31, 29, 32], respectively.
Figure 2: Density plots of the stars with full 6-dimensional kinematic information available from Gaia within 4 kpc of the Solar location in the $x-y$ (top row) and $x-z$ (bottom row) planes. The left column shows all 24,789,061 fully-characterized stars. The middle column shows the 5,811,956 remaining stars after applying the selection criteria described in the text. The right column applies the additional requirement of $|z|>1$ kpc, resulting in 470,702 stars.
Figure 3: Absolute magnitude $M_{G}$ and color $BP-RP$ of all 24,789,061 stars within 4 kpc of the Sun with 6-dimensional kinematic information measured by Gaia. The horizontal dashed white line denotes the magnitude completeness criteria Eq. (2.2). All 5,811,956 stars above the dashed white line are bright enough to be observable for Gaia regardless of position in the 4 kpc sphere. The visible peak in the white box is the red clump. All features in the space of uncorrected $M_{G}$ and $BP-RP$ will appear to be smeared towards the bottom-right of this figure, as dust extinction both dims and reddens stars. While a fraction of red clump stars are smeared beyond the magnitude cut threshold, all of these stars reside in the disk and will not significantly influence solutions to the CBE in the stellar halo.
@@ 2.1 Tracer Population Selection
We first remove from our analysis dataset stars without spectroscopic or photometric magnitudes, as well as stars with large relative parallax errors ($3\sigma_{\varpi}>\varpi$). [^1] This parallax error selection for the dataset quality control has the potential to bias the learned stellar number density, as further stars tend to have larger relative parallax errors. Nevertheless, the fraction of stars removed solely due to this selection criterion is at most 1.5% within the fiducial volume of our analysis, and the statistical uncertainty of flow-learned number density is about 1% around the Sun and about 4% on the boundary of the fiducial volume; hence, the bias is insignificant. We discuss the effect of the relative parallax error selection in Appendix B. The explicit correction of this subleading bias in number density is beyond the scope of this paper, and we will leave it for future studies. Of the 31,532,490 stars in Gaia DR3 with full 6D kinematic solutions available within 10 kpc, 29,855,114 stars remain after these cuts. We further remove 13 abnormally fast stars, with speeds higher than 1000 km/s relative to the Galactic rest frame. Within 4 kpc of the Sun, 24,789,061 stars remain, their position-space distributions are shown in the left column of Figure 2, and their color-magnitude diagram ($BP-RP$ versus $M_{G}$) is shown in Figure 3.
@@
This selection criterion is shown as the white dashed line in Figure 3. Applying this selection removes 63.6% of the stars in the sample with distance $<10$ kpc, leaving $10,876,430$. Note that we place this selection criterion on the magnitude without extinction-correction as the analysis of this paper will focus on the regions with less dust extinction. This approach also helps mitigate issues arising from the non-uniform availability of dust extinction corrections from Gaia across the 4 kpc sphere centered on the Sun. As a result of this non-uniformity, correcting for extinction induces additional spurious position-dependent suppression factors in the derived phase space density, depending on the availability of this information. In the future, these corrections may be more uniformly available, and future versions of Gaia itself and combined catalogs like StarHorse [33] would likely improve this analysis.
@@
Within the dust-filled disk, significant density variations remain after the application of Eq. (2.2). This is most obviously seen as a triangular wedge of apparent low stellar density towards the Galactic center at $|z|\sim 0$, but striations can also be seen in other directions within the disk. These features are primarily in the disk ($|z|\lesssim 1$ kpc) and largely trace known dust features, as can be seen in the left-hand column Figure 5, where we overlay extinction maps of dust [40] with the observed number density of stars after the magnitude selection criterion has been applied. For this comparison, we stitch together two three-dimensional maps: bayestar19 [41] covering declination $>30^{\circ}$, and marshall [42] covering $b\in[-10^{\circ},10^{\circ}]$ and $\ell\in[-100^{\circ},100^{\circ}]$. For the position not covered by either 3D dust maps, we use a 2D dust map SFD [43] for distances $d>1.0\;\mathrm{kpc}$. The regions of low stellar density match the known dust clouds, including the Serpens-Aquila rift, the Rho Ophiuchi cloud complex, Lupus, the Dark River, Pipe Nebula, the Northern Coalsack, and the Vela molecular ridge. In the right column of Figure 2, we show the stars after the magnitude selection Eq. (2.2) with the additional requirement of $|z|>1$ kpc. Neither the effects of crowding nor dust are visible outside the disk.
@@ 3.1 Phase Space Densities
To determine the gravitational accelerations $-\vec{\nabla}\Phi$ from the collisionless Boltzmann Equation applied to a stellar population, we must first determine the phase space density $f(\vec{x},\vec{v})$ for that population. We accomplish this using normalizing flows, an unsupervised machine-learning technique for density estimation. Normalizing flows are based on invertible transformations of a simple base distribution (such as standard normal distribution) into a more complicated distribution. As long as the transformation is expressive enough, normalizing flows are able to model a variety of distributions so that this model can be used as a free-form density estimator. This expressivity is generally achieved using neural networks with bijective constraints.
@@ 3.2 Accelerations
$a_{x}$ $(10^{-10}\mathrm{m/s^{2}})$ $-2.32\pm 0.16$ $-1.94\pm 0.22$
$a_{y}$ $(10^{-10}\mathrm{m/s^{2}})$ $\phantom{-}0.04\pm 0.16$ $\phantom{-}0.08\pm 0.08$
$a_{z}$ $(10^{-10}\mathrm{m/s^{2}})$ $-0.14\pm 0.19$ $-0.06\pm 0.08$
$|\vec{a}|$ $(10^{-10}\mathrm{m/s^{2}})$ $\phantom{-}2.32\pm 0.16$ $\phantom{-}1.94\pm 0.22$
@@
Figure 9: Accelerations $a_{r}$ (top row), $a_{\theta}$ (middle row), and $a_{\phi}$ (bottom row) as a function of spherical radius $r$ (left column), polar arclength $s=r_{\odot}\times(\pi/2-\theta)$ (middle column), and azimuthal arclength $(r_{\odot}^{2}-z^{2})^{1/2}\times\phi$ (right column). Radial and azimuthal measurements are taken along lines off-set from the Galactic midplane, at $z=1.5$ kpc (yellow lines, labeled “North”) and $z=-1.5$ kpc (purple lines, labeled “South”). See Figure 1 for the orientation of these measurement lines within the Galaxy. The inner dark bands and outer light bands denote $1\sigma$ and $2\sigma$ uncertainties respectively. Subpanels in the top and middle rows show the fractional difference between the North and South measurements (appropriately mirrored for $a_{\theta}$). Subpanels in the bottom row show the ratio of $a_{\phi}$ to the magnitude of $\vec{a}$.
In Table 1, we show the estimated acceleration at the Solar location, averaged over a ball centered at the Sun and with a radius 100 pc. [^2] Recall that our data has a small void at the Solar location with radius $\sim 50$ pc, due to the Gaia lower limit on the apparent magnitude. Although our normalizing flows have smoothed out this region, we apply this averaging in order to suppress any artifacts from the interpolation. The measured radial acceleration at the Solar location is $(1.94\pm 0.22)\times 10^{-10}\;\mathrm{m/s^{2}}$; the other components are negligible in comparison. The error includes measurement and statistical errors estimated by resampling and bootstrapping, the variation from multiple independent trainings of the MAFs, and the standard deviation of the estimated acceleration within the 100 pc ball.
@@ 3.4 Mass Densities
[…]
Figure 12: Our averaged measurement of the dark matter density at the Solar radius (top line), compared to recent measurements [13, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 68, 80, 81, 82, 93, 83, 84, 85, 86, 87, 12, 88, 89, 66, 90, 15, 10, 14, 11, 91, 67, 92] of the density of dark matter at or near the Solar location $\rho_{\rm{DM},\odot}$. See [94] for a detailed review of most of these measurements and their techniques. Our measurement of $\overline{\rho}_{\rm{DM}}(r=r_{\odot})=1.38\pm 0.13\times 10^{-2}\,M_{\odot}/{\rm pc}^{3}=0.47\pm 0.05\,{\rm GeV/cm}^{3}$ is consistent with the existing literature of measurements of $\rho_{\rm{DM},\odot}$.
@@
$\overline{\rho}_{\rm DM}(r=r_{\odot})$ $1.18\pm 0.14$ $0.47\pm 0.05\phantom{0}$ $1.38$
@@
with free parameters $\rho_{0}$, $r_{s}$, and $\beta$ (in the standard NFW, $\beta=1$). In performing our fit, we adopted a truncated Gaussian prior for $\beta$ centered at $2$ with a width of 2. We adopt a flat prior for $r_{s}$ within $[0,20]$, and a truncated Gaussian prior centered at $40\times 10^{-2}M_{\odot}/{\rm pc}^{3}$ with a width of $80\times 10^{-2}M_{\odot}/{\rm pc}^{3}$ for $\rho_{0}$ in the range $[0,200]\times 10^{-2}\,M_{\odot}/{\rm pc}^{3}$. This choice of priors restricts the model space to a physically realistic domain: the Gaussian prior for $\rho_{0}$ prevents arbitrarily large density parameters, leading to extremely small values of the scale radius $r_{s}$.
@@ 4 Discussion
We report the acceleration at the Solar location, with $a_{r}=(1.94\pm 0.22)\times 10^{-10}$ m/s2. This is within $2\sigma$ of recent measurements of the acceleration using quasars [65]. Deviations from axisymmetry and the assumptions of equilibrium are found when comparing the measurements in the Galactic North and South. Though at this time the statistical and systematic errors on our MAF-derived accelerations are too large to draw confident conclusions about departures from equilibrium, our method allows us to test assumptions of equilibrium and symmetry in a way that has not been previously possible.

### v3

@@ 1 Introduction
Assuming that the phase space density of the stars is in equilibrium $\partial f/\partial t=0$, the 3D acceleration field $\partial\Phi/\partial x_{i}$ can be derived from knowledge of the stellar phase space density $f$ today. A further derivative of $\Phi$ then gives the total mass density $\rho$; the dark matter density can then be calculated assuming knowledge of the baryonic components.
@@ 2 Gaia DR3
In this analysis, we do not use other spectral information (such as BPRPBP-RP color).
Figure 1: Schematic representation of the Solar location (red dot) relative to the Galactic Center (black dot). The 4 kpc observation volume is shown as a transparent grey sphere. The three coordinate systems used in this work are shown: the Galactocentric Cartesian coordinates $(x,y,z)$, the spherical coordinates $(r,\theta,\phi)$, and the cylindrical coordinates $(R,\phi,z)$. The lines through the observational volume with low dust extinction along which we measure accelerations and mass densities are shown in color. In orange, we show two lines at $z=+1.5$ kpc, one varying $r$ and another $\phi$. The two corresponding lines at z=1.5z=-1.5 kpc are shown in purple. The line parameterized by polar arclength $s=r_{\odot}\times(\pi/2-\theta)$ passing through the Solar location is shown in green.
@@
    Our Cartesian coordinate system places the xyx-y plane in the Galactic disk, with $x=0$ at the Galactic center and the Sun along the +x+x axis. The +z+z axis (perpendicular to the disk) is oriented so that the net rotation of the disk stars in the y-y direction.
@@
    The cylindrical coordinate system $(R,\phi,z)$ uses the same $z$ axis as the Galactocentric Cartesian coordinates. The Sun is located at $\phi=0$, with positive $\phi$ increasing towards the +y+y axis.
@@
    The spherical coordinate system $(r,\theta,\phi)$ has the same $\phi$ angle as the cylindrical coordinates, and measures $\theta$ relative to +z+z axis, with the Galactic disk at $\theta=\pi/2$. We define the polar arc length above or below the Galactic disk at the Solar radius (along the $\theta$ direction) as $s\equiv r_{\odot}\times(\pi/2-\theta)$, with $r_{\odot}=8.122$ kpc.
From the fundamental kinematic properties (parallax, angular position, etc.) measured by Gaia, the positions and velocities in these three coordinate systems can be obtained, using the parallax to calculate distance and assuming a Galactocentric Solar position and velocity of (8.122, 0.0, 0.0208)(8.122,\,0.0,\,0.0208) kpc [29, 30] and (12.9,245.6, 7.78)(-12.9,\,-245.6,\,7.78) km/s [31, 29, 32], respectively.
Figure 2: Density plots of the stars with full 6-dimensional kinematic information available from Gaia within 4 kpc of the Solar location in the xyx-y (top row) and xzx-z (bottom row) planes. The left column shows all 24,789,061 fully-characterized stars. The middle column shows the 5,811,956 remaining stars after applying the selection criteria described in the text. The right column applies the additional requirement of $|z|>1$ kpc, resulting in 470,702 stars.
Figure 3: Absolute magnitude $M_{G}$ and color BPRPBP-RP of all 24,789,061 stars within 4 kpc of the Sun with 6-dimensional kinematic information measured by Gaia. The horizontal dashed white line denotes the magnitude completeness criteria Eq. (2.2). All 5,811,956 stars above the dashed white line are bright enough to be observable for Gaia regardless of position in the 4 kpc sphere. The visible peak in the white box is the red clump. All features in the space of uncorrected $M_{G}$ and BPRPBP-RP will appear to be smeared towards the bottom-right of this figure, as dust extinction both dims and reddens stars. While a fraction of red clump stars are smeared beyond the magnitude cut threshold, all of these stars reside in the disk and will not significantly influence solutions to the CBE in the stellar halo.
@@ 2.1 Tracer Population Selection
We first remove from our analysis dataset stars without spectroscopic or photometric magnitudes, as well as stars with large relative parallax errors ($3\sigma_{\varpi}>\varpi$). [^1] This parallax error selection for the dataset quality control has the potential to bias the learned stellar number density, as further stars tend to have larger relative parallax errors. Nevertheless, the fraction of stars removed solely due to this selection criterion is at most 1.5% within the fiducial volume of our analysis, and the statistical uncertainty of flow-learned number density is about 1% around the Sun and about 4% on the boundary of the fiducial volume; hence, the bias is insignificant. We discuss the effect of the relative parallax error selection in Appendix B. The explicit correction of this subleading bias in number density is beyond the scope of this paper, and we will leave it for future studies. Of the 31,532,490 stars in Gaia DR3 with full 6D kinematic solutions available within 10 kpc, 29,855,114 stars remain after these cuts. We further remove 13 abnormally fast stars, with speeds higher than 1000 km/s relative to the Galactic rest frame. Within 4 kpc of the Sun, 24,789,061 stars remain, their position-space distributions are shown in the left column of Figure 2, and their color-magnitude diagram (BPRPBP-RP versus $M_{G}$) is shown in Figure 3.
@@
This selection criterion is shown as the white dashed line in Figure 3. Applying this selection removes 63.6% of the stars in the sample with distance <10<10 kpc, leaving $10,876,430$. Note that we place this selection criterion on the magnitude without extinction-correction as the analysis of this paper will focus on the regions with less dust extinction. This approach also helps mitigate issues arising from the non-uniform availability of dust extinction corrections from Gaia across the 4 kpc sphere centered on the Sun. As a result of this non-uniformity, correcting for extinction induces additional spurious position-dependent suppression factors in the derived phase space density, depending on the availability of this information. In the future, these corrections may be more uniformly available, and future versions of Gaia itself and combined catalogs like StarHorse [33] would likely improve this analysis.
@@
Within the dust-filled disk, significant density variations remain after the application of Eq. (2.2). This is most obviously seen as a triangular wedge of apparent low stellar density towards the Galactic center at $|z|\sim 0$, but striations can also be seen in other directions within the disk. These features are primarily in the disk ($|z|\lesssim 1$ kpc) and largely trace known dust features, as can be seen in the left-hand column Figure 5, where we overlay extinction maps of dust [40] with the observed number density of stars after the magnitude selection criterion has been applied. For this comparison, we stitch together two three-dimensional maps: bayestar19 [41] covering declination $30^{\circ}$, and marshall [42] covering $b\in[-10^{\circ},10^{\circ}]$ and $\ell\in[-100^{\circ},100^{\circ}]$. For the position not covered by either 3D dust maps, we use a 2D dust map SFD [43] for distances $d>1.0\;\mathrm{kpc}$. The regions of low stellar density match the known dust clouds, including the Serpens-Aquila rift, the Rho Ophiuchi cloud complex, Lupus, the Dark River, Pipe Nebula, the Northern Coalsack, and the Vela molecular ridge. In the right column of Figure 2, we show the stars after the magnitude selection Eq. (2.2) with the additional requirement of $|z|>1$ kpc. Neither the effects of crowding nor dust are visible outside the disk.
@@ 3.1 Phase Space Densities
To determine the gravitational accelerations $\vec{\nabla}\Phi$ from the collisionless Boltzmann Equation applied to a stellar population, we must first determine the phase space density $f(\vec{x},\vec{v})$ for that population. We accomplish this using normalizing flows, an unsupervised machine-learning technique for density estimation. Normalizing flows are based on invertible transformations of a simple base distribution (such as standard normal distribution) into a more complicated distribution. As long as the transformation is expressive enough, normalizing flows are able to model a variety of distributions so that this model can be used as a free-form density estimator. This expressivity is generally achieved using neural networks with bijective constraints.
@@ 3.2 Accelerations
$a_{x}$ $(10^{-10}\mathrm{m/s^{2}})$ $2.32\pm 0.16$ $1.93\pm 0.08$
$a_{y}$ $(10^{-10}\mathrm{m/s^{2}})$ $\phantom{-}0.04\pm 0.16$ $\phantom{-}0.08\pm 0.05$
$a_{z}$ $(10^{-10}\mathrm{m/s^{2}})$ $14\pm 0.19 0.0$$05\pm 0.06$
$|\vec{a}|$ $(10^{-10}\mathrm{m/s^{2}})$ $\phantom{-}2.32\pm 0.16$ $\phantom{-}1.93\pm 0.08$
@@
Figure 9: Accelerations $a_{r}$ (top row), $a_{\theta}$ (middle row), and $a_{\phi}$ (bottom row) as a function of spherical radius $r$ (left column), polar arclength $s=r_{\odot}\times(\pi/2-\theta)$ (middle column), and azimuthal arclength $(r_{\odot}^{2}-z^{2})^{1/2}\times\phi$ (right column). Radial and azimuthal measurements are taken along lines off-set from the Galactic midplane, at $z=1.5$ kpc (yellow lines, labeled “North”) and z=1.5z=-1.5 kpc (purple lines, labeled “South”). See Figure 1 for the orientation of these measurement lines within the Galaxy. The inner dark bands and outer light bands denote $1\sigma$ and $2\sigma$ uncertainties respectively. Subpanels in the top and middle rows show the fractional difference between the North and South measurements (appropriately mirrored for $a_{\theta}$). Subpanels in the bottom row show the ratio of $a_{\phi}$ to the magnitude of $\vec{a}$.
In Table 1, we show the estimated acceleration at the Solar location, averaged over a ball centered at the Sun and with a radius 100 pc. [^2] Recall that our data has a small void at the Solar location with radius $\sim 50$ pc, due to the Gaia lower limit on the apparent magnitude. Although our normalizing flows have smoothed out this region, we apply this averaging in order to suppress any artifacts from the interpolation. The measured radial acceleration at the Solar location is $(-1.93\pm 0.08)\times 10^{-10}\;\mathrm{m/s^{2}}$; the other components are negligible in comparison. The error includes measurement and statistical errors estimated by resampling and bootstrapping, the variation from multiple independent trainings of the MAFs, and the standard deviation of the estimated acceleration within the 100 pc ball.
@@ 3.4 Mass Densities
Refer to captionThis workCasagrande, (2020) [69]Pato, et al., (2015) [70]Huang, et al., (2016) [71]Benito, et al., (2019) [72]Karukes, et al., (2019) [73]Lin, et al., (2019) [74]de Salas, et al., (2019) [75]Ablimit, et al., (2020) [76]Benito, et al., (2020) [77]Sofue, (2020) [78]Zhou, et al., (2022) [79]Ou, et al., (2023) [68]McMillan, (2017) [80]Cautun, et al., (2020) [81]Bienyame, et al., (2014) [82]Piffl, et al., (2014) [83]Binney, et al., (2015) [84]Cole, et al., (2017) [85]Wegg, et al., (2019) [86]Hattori, et al., (2020) [87]Nitschai, et al., (2020) [12]Nitschai, et al., (2021) [13]Widmark, et al., (2021) [88]Guo, et al., (2022) [89]McKee, et al., (2015) [66]Xia, et al., (2016) [90]Hagen, et al., (2018) [15]Sivertsson, et al., (2018) [10]Guo, et al., (2020) [14]Salomon, et al., (2020) (North) [11]Salomon, et al., (2020) (South) [11]Wardana, et al., (2020) [91]Schutz, et al., (2018) [67]Buch, et al., (2019) [92]
Figure 12: Our averaged measurement of the dark matter density at the Solar radius (top line), compared to recent measurements [13, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 68, 80, 81, 82, 93, 83, 84, 85, 86, 87, 12, 88, 89, 66, 90, 15, 10, 14, 11, 91, 67, 92] of the density of dark matter at or near the Solar location $\rho_{\rm{DM},\odot}$. See [94] for a detailed review of most of these measurements and their techniques. Our measurement of $\overline{\rho}_{\rm{DM}}(r=r_{\odot})=1.23\pm 0.13\times 10^{-2}\,M_{\odot}/{\rm pc}^{3}=0.47\pm 0.05\,{\rm GeV/cm}^{3}$ is consistent with the existing literature of measurements of $\rho_{\rm{DM},\odot}$.
@@
$\overline{\rho}_{\rm DM}(r=r_{\odot})$ $1.23\pm 0.13$ $0.47\pm 0.05\phantom{0}$ $1.38$
@@
with free parameters $\rho_{0}$, $r_{s}$, and $\beta$ (in the standard NFW, $\beta=1$). In performing our fit, we adopted a truncated Gaussian prior for $\beta$ centered at $2$ with a width of 2. We adopt a flat prior for $r_{s}$ within $[0,20]$, and a truncated Gaussian prior centered at $40\times 10^{-2}M_{\odot}/{\rm pc}^{3}$ with a width of $80\times 10^{-2}M_{\odot}/{\rm pc}^{3}$ for $\rho_{0}$ in the range $0,200]\times 10^{-2}\,M_{\odot}/{\rm pc}^{3}$. This choice of priors restricts the model space to a physically realistic domain: the Gaussian prior for $\rho_{0}$ prevents arbitrarily large density parameters, leading to extremely small values of the scale radius $r_{s}$.
@@ 4 Discussion
We report the acceleration at the Solar location, with $a_{r}=(-1.93\pm 0.08)\times 10^{-10}$ m/s2. This is almost within $2\sigma$ of recent measurements of the acceleration using quasars [65]. Deviations from axisymmetry and the assumptions of equilibrium are found when comparing the measurements in the Galactic North and South. Though at this time the statistical and systematic errors on our MAF-derived accelerations are too large to draw confident conclusions about departures from equilibrium, our method allows us to test assumptions of equilibrium and symmetry in a way that has not been previously possible.